# Retinomorphic Machine Vision in a Network Laser

### v3

@@ Methods
Supplementary Information
Optical setup schematic
Figure S1 shows a visual schematic of the experimental setup and optical path.
Figure S1: Experimental optical setup.
Network laser image and spectral location of feature-detecting modes
Figure S2 shows details of the InP network laser structure, spatial localisation of lasing modes and the hyperpectral regions used for feature detection. Figure S2a) shows an optical microscope image of the network used in this study. Figures S2 b) and c) show the spatial profile of the light emitted by the network in response to a spatially-structured pump illumination (dark regions of the network receive no pump illumination. Figure S2 b) shows all light emitted by the networks, figure S2 c) shows the spatial localisation of different wavelength lasing modes, with the colour-coding corresponding to the highlighted regions on the hyperspectra shown in figure S2 d). The colour corresponds to the scheme used for the mode feature maps throughout this work, eg. the red feature maps are formed via the amplitude of the hyperspectral region highlighted in red (labelled ‘5’). Figure S2 e) shows the spectral position of the feature-detecting mode formed by averaging over all spatial channels.
Figure S2: a) Optical microscope image of the InP network laser used in this study.
b) Optical image of light emitted by network under diagonal-edge illumination (dark regions of the network receive no pump illumination.
c) Optical image from b) with emitted light by the network colour-coded corresponding to the wavelength of the emitted light, with the colours corresponding to the labelled hyperspectral regions in d). Different regions of the network are seen to emit light at different wavelengths.
d) Network hyperspectral output recorded for the pump illumination pattern shown in b,c). Nine hyperspectral regions are labelled, colour-coded in correspondence with the detected feature maps used throughout this study indicating the wavelength and spatial position of the spectral regions used for feature detection. Notably, feature detecting regions are largely positioned around the edges of the main central region of hyperspectral intensity.
e) Position of the tenth feature-detecting spectral region, which is formed by averaging over all spatial channels and taking the amplitude within the indicated grey region.
NetSALT simulated feature detection - additional feature maps and spatial mode profiles
Figure S3 shows an expanded set of 28 netSALT simulated modes, spatial mode profiles and corresponding feature maps for the CL test image. A broad range of detected image features are observed.
Figure S4 shows the full set of all 172 netSALT simulated mode feature maps. Many detected features are observed, alongside modes which only respond to very specific image regions - often lasing for only a handful of pixels of the test image eg modes 62, 157 and 109. Some modes return an output close to the input image, sometimes with enhanced lasing amplitude along a specific image feature, eg modes 2, 6, 9 and 35.
Figure S3: Spectrally-multiplexed feature-maps and corresponding spatial mode profiles simulated via netSALT.
Figure S4: Feature maps from all 172 netSALT simulated modes.
NetSALT simulated feature detection for a regular hexagonal network
Figure S5 shows netSALT simulated lasing spectra and feature detection for a network with a regular hexagonal graph topology. The regular network shows some ability to detect features, seen in figures S5 c,d) where the system can produce different outputs between full illumination, left/right edge illumination and top/bottom edge illumination - but the regular network can not discern the difference between left and right edges or top and bottom edges. Looking at the full range of detected feature maps in S5d), a much smaller range of features are detected relative to the irregular network.
Figure S5: netSALT simulations of a regular hexagonal network. (a) Regular hexagonal graph topology, 50 micron diameter graph. (b) Comparison of network spectra under full illumination between the regular hexagonal network and the irregular network examined in fig 2. (c) Comparison of regular hexagonal network lasing spectra under a variety of illumination patterns, full illumination and left, right, bottom and top edge illumination. The spectral response can differentiate between full illumination, left/right edge illumination and top/bottom edge illumination - but not between left and right edges, or top and bottom edges. The irregular network provides more sensitivity across a greater range of image features than the regular network. (d) Mode feature maps for all the active regular graph modes. Far fewer features are detected than the irregular graph.
Experimental feature detection - additional feature maps and test images
Figures S6, S7 and S8 show experimentally measured feature maps for 10 lasing modes for a range of test images to further examine the feature detection functionality.
Figures S6 and S7 show feature maps for modes 0-4 and 5-9 respectively in response to four test images: positive and negative tone circles and stars. The feature maps for these test images give a strong visual representation of the range of image features which each mode is sensitive to. Each mode responds to a distribution of similar edges, e.g. some finite arc on the circle or a range of adjacent spikes on the star. This demonstrates that the modes don’t only respond to one very specific illumination pattern, but rather a distribution of similar features. This is exactly the behaviour desired for feature detection functionality, eg mode 3 responds strongly across a range of ‘left’ image edges, which is far more useful than only responding to a perfect vertical left edge.
Figure S8 shows the response of modes 0-9 for a range of image edges (input images in left-most column) which are gradually rotated. This provides a good visual indication of the range of image features detected by each mode, and also shows that the ten mode provide feature detection response across all angles of image edge between them.
It is important to note that while in netSALT we can spectrally isolate the lasing response of specific modes, even if multiple modes may occupy overlapping spectral regions - this is not possible in experiment. As a result, the selected experimental hyperspectral regions and corresponding feature maps may be contributed to by more than one lasing mode.
Figures S10, S11, S12, S13 and S14 show feature maps for all ten modes across the cameraman, CL and street test images, and selected MNIST-digits and Fashion-MNIST images respectively.
Figure S6: Experimental feature maps and learned equivalent kernels for modes 0-5 on four test images, of black and white circles and stars. Each mode has a kernel displayed which was learned by comparing input images to experimentally measured feature maps, and learning the effective kernel via linear regression.
Figure S7: Experimental feature maps and learned equivalent kernels for modes 6-9 (mode 9 integrates over all spatial channels, hence is labelled lambda as it contains no spatial information). on four test images, of black and white circles and stars. Each mode has a kernel displayed which was learned by comparing input images to experimentally measured feature maps, and learning the effective kernel via linear regression.
Figure S8: Experimental feature maps for modes 0-9 for input images (left-most column) of a range of rotated edges, and two white noise image inputs (bottom two rows). Each mode has two columns, an analogue-varying output produced by the mode-amplitude (left column for each mode, labelled with mode number) and a binarised black and white output produced by normalising the analogue output to 0 or maximum pixel intensity (right column for each mode, labelled BW.)
Least-squares fitting of kernel filters
As described in the main text and shown in figure 2i), we perform linear least-squares fitting between input images and feature detection output to fit effective linear kernel filters for the experimental modes performing feature detection. Results of this process are shown in figure S9. The fitting process uses a least-squares method to find the linear kernel filter weights which best reproduce the filter output (e.g. the lasing mode amplitude for our experimental case) for a given set of input images and filter outputs. As our input image set, we use the 65,536 ($2^{16}$) possible $4\times 4$ black and white pixel images.
We plot the results of the fitting process for our ten experimental feature-detecting modes in the blue left-hand column of figure S9, titled ‘experimental fits’. For the 65,536 input images we show the target output (eg. the experimental lasing mode amplitude) and residual fitting errors in the blue and orange traces respectively in the column titled ‘mode output’ next to the fitted kernels. Substantial residual errors are observed, with R2 values of 0.628-0.946. This suggests that the linear fitting process may be unable to fully capture the feature detecting dynamics of the experimental system.
To test this assumption, we also performed the same fitting process on linear software kernel filters, initialising ten $4\times 4$ filters with randomly generated weights. The random kernels are plotted in the left-hand orange column titled ‘random kernels’. The results of the linear fitting process for these kernels are shown in the adjacent columns titled ‘fit’ (the kernel weights) and ‘convolved with 4$\times$4 images’ (target and residual errors). Here, perfect R2 scores of 1.0 are observed, as expected when performing a linear fit on a linear process. To test whether the poorer-quality linear fits to our experimental data are a result of experimental noise, we performed a further test where we added Gaussian noise to the software filter output. We used a noise amplitude of 10% relative to the maximum software kernel weights. Results of this test are shown in the two right-most columns. Lower $R^{2}$ of 0.978-0.988 are observed relative to the no-noise case, but still substantially higher than the fits to the experimental data.
Figure S9: Least-squares fitting of kernel filters. Left-hand blue column shows the results of the fitting process for the ten experimental feature detecting modes. R2 values from 0.628-0.946 are found, suggesting aspects of the experimental feature detection functionality are not fully captured by the linear fitting. The orange and blue mode output traces show the target value to be fitted (blue) and the fitting error residual (orange) for the 65,536 ($2^{16}$) possible $4\times 4$ pixel images. The central orange column shows the same process performed for randomly generated linear software kernel filters, showing perfect R2 scores of 1.0. The right-hand column shows the same process performed where Gaussian noise of 10% amplitude relative to the maximum kernel weight is added, with reduced but still high R2 scores of 0.978-0.988.
Extracting effective mode kernel filters from experimental data
To compare the feature detection functionality of our system against the convolutional kernel filter approach used in software CNNs, we extracted effective kernel filters for our lasing modes. This was done by performing least-squares fitting between all possible 4$\times$4 pixel black and white input images (65,536 images) and experimentally measured mode amplitudes for each of the possible input images.
The results of this are shown in figure 2i) alongside $R^{2}$ values for each extracted effective kernel filter. When this fitting process is performed for an arbitrary linear kernel generated in software, the fitting process performs perfectly with $R^{2}$ values of 1.0. For our experimental modes, the fitting process returns $R^{2}$ values between 0.76-0.95, suggesting that aspects of the physical feature detection process cannot be captured by linear matrix kernel filters.
Figure S10: Experimental feature maps of the ‘cameraman’ test image for modes 0-9.
Figure S11: Experimental feature maps of ‘CL’ test image for modes 0-9.
Figure S12: Experimental feature maps of ‘street’ test image for modes 0-9.
Figure S13: Experimental feature maps of ten MNIST digits test images, one per digit class, for modes 0-9.
Figure S14: Experimental feature maps of ten Fashion-MNIST test images, one per object class, for modes 0-9.
Quasi-isomorphic software benchmarks
To compare our system against a software network with similar architecture, we constructed a network with quasi-isomorphic structure to our experimental system. The quasi-isomorphic software benchmark comprises an initial convolutional layer with 10 4$\times$4 convolutional filters, followed by a hidden layer of 30,000 neurons with randomised weights and ReLu nonlinear activation, with a final logistic regression layer. We compared our experimental system against two versions of this quasi-isomorphic software network: one with randomised kernel filter weights, and one with kernel filter weights obtained via fits to our experimental data as shown in figure 2f) and S6, S7.
Figures S15 and S16 show the classification results of these tests for the MNIST digits and Fashion-MNIST tasks respectively. The experimental photonic system outperforms the quasi-isomorphic benchmark in both cases, with the software benchmark using the fitted experimental kernel weights performing better.
Figure S15: MNIST digits classification results for the experimental photonic system and two versions of the quasi-isomorphic software benchmark, one with random kernel filter weights and one with kernel filter weights obtained via fits to our experimental data.
Figure S16: Fashion MNIST classification results for the experimental photonic system and two versions of the quasi-isomorphic software benchmark, one with random kernel filter weights and one with kernel filter weights obtained via fits to our experimental data.
Speed, energy and scalability
For the measurements taken in this study, each image illumination takes approximately 300 ms. As we were preparing the paper, the scheme has already been improved to run at 3 ms per image by increasing the laser repetition rate from 1 kHz to 100 kHz and decreasing spectrometer integration time, and there is a clear route to future scalability through further engineering optimisation of illumination and detection - the physical lasing dynamics of the system take 100 ps illumination. We are integrating multiple parallel processing steps in each illumination: 10 parallel feature detection processes, each with a complex nonlinear activation.
The 300 ms time per image is largely due to the fact that A) we are not using a high-speed spectrometer and are saving the data on a low-end PC without high-speed digital electronics such as a fast capture card or FPGA, and B) we operated our pump laser at a relatively low repetition rate of 1 kHz for these measurements, and integrated over 300 pulses per image to improve signal-to-noise. As mentioned above, while preparing this manuscript we have improved our experimental implementation and increased image throughput by a factor of 100$\times$.
Commercially available Si or InGaAs photodiode arrays (e.g. Hamamatsu) run from $10^{4}-10^{6}$ Hz and are readily integrable with our scheme. There is ample headroom for increasing our throughput via simple component upgrades and the 100 ps timescale of the key physical dynamics is attractive.
Currently we are using pump illuminations that supply 43 nJ per illumination to the InP network sample (experimentally measured via power meter). Energy consumption could be further reduced by optimisation of our network topology and geometry including lowering the system lasing threshold, for instance we found during the final stages of this work that using networks with 450 nm wide waveguides reduces the lasing threshold by around 50% relative to our existing waveguide width of 350 nm[29], substantially reducing the required energy per illumination. There is scope for further reducing the lasing threshold through similar future device optimisations.

### v4

@@ Methods
EfficientNetV2-B0 Transfer Learning Model
This model uses transfer learning using the EfficientNetV2-B0 architecture with an added classification head MLP. EfficientNetV2 is a relatively modern (2019[40] original EfficientNet model, 2021[41] for the EfficientNetV2 employed here) large convolutional neural network, frequently deployed in industrial and commercial learning applications and applied to real world problems. EfficientNetV2-B0 has 3040 convolutional filters, far more powerful than the other CNNs considered here. As is common in learning applications without sufficiently large training sets to fully optimise network parameters from scratch, pretrained weights learned from training on the ImageNet dataset are used to initialise the network and set up a broad range of feature extracting kernels before the network is optimised on the training set of an arbitrary desired task.
The EfficientNetV2-B0 base model is followed by a 512 $\times$ 128 neuron multilayer perceptron classification head to aid repurposing of the initialised ImageNet weights to arbitrary tasks, comprising: global average pooling, two fully connected (dense) layers with ReLU activations, and a final softmax classification layer.
  • EfficientNetV2-B0 (pretrained on ImageNet) with 512 $\times$ 128 MLP head
    • Stem Convolution Layer:
        Filters: 32
        Kernel Size: $3\times 3$
        Stride: 2
        Activation: Swish
    • Fused-MBConv Blocks:
        Stage 1: 1 block, 16 filters, expansion: 1, kernel size: $3\times 3$
        Stage 2: 2 blocks, 24 filters, expansion: 4, kernel size: $3\times 3$
        Stage 3: 2 blocks, 32 filters, expansion: 4, kernel size: $3\times 3$
    • MBConv Blocks:
        Stage 4: 3 blocks, 48 filters, expansion: 4, kernel size: $3\times 3$
        Stage 5: 5 blocks, 64 filters, expansion: 6, kernel size: $3\times 3$
        Stage 6: 8 blocks, 128 filters, expansion: 6, kernel size: $3\times 3$
        Stage 7: 1 block, 160 filters, expansion: 6, kernel size: $3\times 3$
    • Head Convolution Layer:
        Filters: 1280
        Kernel Size: $1\times 1$
        Activation: Swish
    • Global Average Pooling:
        Reduces spatial dimensions to a single 1280-dimensional feature vector
    • Dense Layer 1:
        Neurons: 512
        Activation: ReLU
    • Dense Layer 2:
        Neurons: 128
        Activation: ReLU
Number of Trainable Parameters:
EfficientNetV2-B0 Base: $\approx 7.2\text{ million parameters}$
Dense Layer 1: $(1280+1)\times 512=655,872$
Dense Layer 2: $(512+1)\times 128=65,664$
Output Layer: $(128+1)\times 10=1,290$
Total: $\approx 7,922,826\text{ parameters}$